OpenStudy (anonymous):

Given g(x)=1.1845-log x determine for what value of x that g(x)=1

OpenStudy (danjs):

g(x) = 1 = 1.1845 - log(x)

OpenStudy (danjs):

log(x) = 1.1845 - 1 log(x) = 0.1845

OpenStudy (danjs):

10^log(x) = x = 10^(0.1845) = calculator number

OpenStudy (anonymous):

thanks. Do you know of an instructional site that teaches you how to use your calculator? I am unable to move past the basic keys to solve equations.

OpenStudy (danjs):

should be a button that looks like y^x or x^y so you can solve x =10^0.1845

OpenStudy (danjs):

not sure, you would have to search for your specific calculator instructions.

OpenStudy (anonymous):

I will Thank you so much.